For the given rational function f, find all values of x for which f(x) has the indicated value.
f(x) = 5x + 21/x ; f(x) = -38

Answer by RAY100(1637)

f(x) = -38 = 5x +21/x
.
mult thru by x
.
-38x = 5x^2 +21
.
0=5x^2 +38x +21
.
factor
.
(5x+3)(x+7) =0
.
x= -3/5,,,-7
.
check
(-7),,,-38=5(-7) +21/(-7)
.
-38 = -35 -3,,,,,,ok
.
(-3/5),,,,,-38 =5(-3/5) + 21/(-3/5)
.
-38 = -3 -35 = -38,,,,,ok